OpenStudy (anonymous):

find the nth term if the sum of n terms is given by S_n = n^2 + n + 1

hartnn (hartnn):

find sum upto n-1 terms , i.e, fin \[S_{n-1}\] then n'th term = \[S{n}-S_{n-1}\]

OpenStudy (anonymous):

well i've tried that already that doesnt work here, the constant term is causing problems

hartnn (hartnn):

means you have answer and your work doesn't match with answer given ?

OpenStudy (anonymous):

i dont have the answer.

OpenStudy (anonymous):

and the answer which comes out from this method, if you cross check it by putting different values of n, doesnt satisfy.

hartnn (hartnn):

okk...show your work...

OpenStudy (anonymous):

why dont you give it a try, am quite sure the method wont work.

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

Sn = n^2 + n + 1 Sn-1 = (n-1)^2 + n Sn -Sn-1 = an = n^2 +1 - (n-1)^2 now lets check : a1 = 2 but S1 = 3 contradiction ..

hartnn (hartnn):

yes, even i got n'th term =2n is this arithmetic series ?

OpenStudy (anonymous):

i dont know what series is this, my teacher asked me this, she couldnt solve it herself.

OpenStudy (anonymous):

i tried another way also, here is where i got stuck : Sn=n^2 + n +1 =n(n+1) +1 =2 summation(n) + 1 had 1 not been there, answer would have been 2n, but the constant is causing problems.

OpenStudy (anonymous):

only if we knew summation of what sequence/series =1 ..can we figure that out ?

OpenStudy (anonymous):

i mean summation upto n terms of what sequence/series =1 ?

hartnn (hartnn):

i can think of one such infinite series, but i don't think that will be expected here.... 0.5+0.25+0.125+... =1

OpenStudy (anonymous):

n should be taking any values/natural number. yours is a particular case of n->inf

OpenStudy (anonymous):

@ParthKohli @shubhamsrg @mukushla

OpenStudy (anonymous):

ohh,,damn clicked closed where i had to click bump ! >.< am i allowed to reopen ?

OpenStudy (anonymous):

the question* ?

OpenStudy (anonymous):

hmm,,guess i'll reask the question.

OpenStudy (anonymous):

here.. http://openstudy.com/study#/updates/50ee7e29e4b07cd2b649a374